OpenStudy (sarahann):

Find the surface area of a pyramid with area of the base 20, slant height 7, and perimeter of the base 30.

OpenStudy (blurbendy):

S = B + (1/2)Pl, where B is the base area, P is the perimeter of the base, and l is the slant height. Given that the base area is 20, the slant height is 7, and the perimeter of the base is 30, we see that: S = 20 + (1/2)(30)(7) = 125 units^2.